OpenStudy (anonymous):

fix n,k belongs to N and consider p1 and p2, both primes. find how many divisors b= p1^k p2^n has and prove your result. conjecture and provide some justification how many divisors there are for a number that can be represented as the product of m arbitrary primes, each raised to a different power.

ganeshie8 (ganeshie8):

Hey!

OpenStudy (anonymous):

what's up?

ganeshie8 (ganeshie8):

have you figured out the solution ?

OpenStudy (anonymous):

not yet

ganeshie8 (ganeshie8):

tried anything ?

ganeshie8 (ganeshie8):

Let \(b = 2^1*3^2\) is it easy to list all the positive divisors of \(b\) ?

OpenStudy (anonymous):

i don't know how to start

ganeshie8 (ganeshie8):

what are the divisors of \(18\) ?

ganeshie8 (ganeshie8):

does 6 divide 18 ?

OpenStudy (anonymous):

1, 2, 3, 6, 9, 18

OpenStudy (anonymous):

ok how can i know the form in general please?

OpenStudy (anonymous):

i tried some other products and i got 6 divisors each time!

OpenStudy (anonymous):

oh no for 2^3 * 3=24, it has 8 divisors which are 1, 2, 3, 4, 6, 8, 12, 24

OpenStudy (anonymous):

i cannot figure out the formula to know how many divisors?

ganeshie8 (ganeshie8):

Excellent! you're very close to a formula...

ganeshie8 (ganeshie8):

Consider \(b = 2^3*3\)

ganeshie8 (ganeshie8):

The key thing to observe here is that any divisor of \(b\) will be of form \(2^a*3^b\) where \(0\le a\le 3\) and \(0\le b\le 1\)

ganeshie8 (ganeshie8):

for example, a = 2, b = 1 gives \(2^2*3^1 = 12\) as one divisor

ganeshie8 (ganeshie8):

\(0\le a\le 3\) how many choices are there for \(a\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

good \(0\le b\le 1\) how many choices are there for \(b\) ?

ganeshie8 (ganeshie8):

Yes, 4*2 = ?

OpenStudy (anonymous):

8

ganeshie8 (ganeshie8):

so, there are 8 divisors of 2^3*3

ganeshie8 (ganeshie8):

you can choose "2" in "3+1" ways you can choose "3" in "1+1" ways overall, number of divisors is given by (3+1)(1+1)

ganeshie8 (ganeshie8):

lets do another example

ganeshie8 (ganeshie8):

Find the number of divisors of \(2^2*5^2\)

OpenStudy (anonymous):

0≤a≤2 3 choices 0≤b≤2 3 choices 3*3=9

OpenStudy (anonymous):

so the number of divisors is (a+1)*(b+1)

OpenStudy (anonymous):

in general (a+1)*(b+1)...(n+1)

OpenStudy (anonymous):

cool! can you help me please how to prove that ?

OpenStudy (anonymous):

oh i got a question 3^2 * 5=45 has 5 divisors: 1, 3, 5, 15, 45 but if i did the formula 0≤a≤2 3 choices 0≤b≤1 2 choices 3*2=6 not 5 why is it different?

ganeshie8 (ganeshie8):

looks you have missed one divisor

ganeshie8 (ganeshie8):

9 ?

OpenStudy (anonymous):

oh yeah thanks